### v1

# Goal-Oriented Multiple Access Connectivity for Networked Intelligent Systems

Abstract
We design a self-decision goal-oriented multiple access scheme, where sensing agents observe a common event and individually decide to communicate the event’s attributes to the monitoring agents, to satisfy a certain goal. Decisions are based on the usefulness of contents, which are generated under uniform, change- and semantics-aware content acquisition, as well as statistics and contents of other agents. We obtain optimal activation probabilities and threshold criteria for decision-making under all schemes, maximizing a grade of effectiveness metric. Combined with a semantics-aware acquisition scheme, the self-decision scheme offers, on average, $29.52\%$ higher effectiveness, $25.13\%$ fewer drop-offs, and $67.21\%$ fewer transmissions.
I Introduction
Goal-oriented semantic communication holds great promise for realizing resource-efficient and scalable networked intelligent systems [1, 2, 3], where sensing agents observe events and convey the events’ attributes to monitoring agents to achieve a certain goal. Therein, content acquisition, data transmission, and information usage become effective in satisfying the goal when only useful or significant content is considered in the communication lifecycle. The challenge is exacerbated when the medium is shared, and agents have to communicate over multiple access channels. Conventional medium access protocols, such as typed-based ALOHA, grant-based, and grant-free access, are primarily goal- and effectiveness-agnostic, sending all frames regardless of their goal-dependent usefulness and impact at the endpoint. Despite some prior work [4, 5, 6, 7, 8, 9, 10], the problem of goal-oriented multiple access and content provision remains mostly unexplored.
In this letter, we propose a self-decision goal-oriented multiple access scheme, in which each sensing agent decides to speak up or remain silent based on the usefulness of its generated contents and the statistics of the other agents and their generated contents for achieving a specific goal. The design involves computing the optimal activation probabilities of the agents and a decision-making criterion, which maximizes a grade of effectiveness metric that integrates the discrepancy error between actual and reconstructed events, the consumed resources, and the usefulness of perceived contents. Our results reveal that the proposed self-decision scheme combined with semantics-aware acquisition offers higher effectiveness with fewer content transmissions and channel drop-offs.
II System Model
[…]
[scale=0.35]Images/SystemModel.eps
Figure 1: Goal-oriented medium access in a networked intelligent system.
We consider $K$ intelligent sensing agents (ISAs) that observe a common event (information source) and convey their noisy observations to $M$ networked monitoring agents (NMAs) over a multiple access wireless channel prone to errors and universal frequency reuse (see Fig. 1). Communication occurs over independent time-slotted service periods. At each service period, all NMAs query for new observations from the ISAs in the downlink, and the ISAs respond to those queries in the uplink, providing content within the same service period. The evolution of the source $X_{t}$ at the $t$-th service period, with $t\in\mathbb{N}$, belongs to a finite state space. At each state, the event is described via a set of information attributes, defined by $\mathcal{A}=\{\mathcal{X}_{n}~{}|~{}n=1,2,…,N\}$. Here, $\mathcal{X}_{n}\in\{x_{i}^{(n)}~{}|~{}i=1,2,…,I_{n}\}$ is a level-$n$ attribute, which is notated as $\text{AL}_{n}$ and modeled by a discrete-time Markov chain (DTMC) with $I_{n}\geq 2$ states, as shown in Fig. 1. In the DTMC modeling the $\text{AL}_{n}$, we assume $p_{i,i}=q_{n}$, and $p_{i,i^{\prime}}=p_{n}$, $\forall i\neq i^{\prime}$, for $i,i^{\prime}=1,2,...,I_{n}$. Therefore, we have $q_{n}+(I_{n}\!-\!1)p_{n}=1$.
In such model, at the $t$-th service period, the NMAs query for a group of attributes $\widehat{\mathcal{A}}_{t}\subseteq\mathcal{A}$ represented by the sequence $\langle\widehat{\mathcal{A}}_{t}\rangle$, where $|\widehat{\mathcal{A}}_{t}|\leq N$. Either queried attribute is independently communicated in one slot $j$, $j=1,2,…,|\widehat{\mathcal{A}}_{t}|$, with the same order as in $\langle\widehat{\mathcal{A}}_{t}\rangle$. The levels of queried attributes are determined in line with the goal to which the NMAs subscribe. The $k$-th ISA, where $k=1,2,…,K$, can observe only a subset of attributes $\mathcal{A}_{k}\subseteq\mathcal{A}$ and generate content from its noisy observations via an encoder, yielding a codebook shared with the other agents. Moreover, $\mathcal{K}_{n}=\langle k:\mathcal{X}_{n}\in\mathcal{A}_{k}\rangle$ is the sequence of ISAs that can observe the $n$-th attribute. To estimate the event, the NMAs need to jointly decode all queried attributes and construct $\widehat{X}_{t}$ at the $t$-th service period. An attribute is correctly perceived if at least $\lceil\frac{M+1}{2}\rceil$ NMAs successfully decode it.
II-A Goal-Oriented Content Provision
At the time slot order of the $n$-th queried attribute, the $k$-th ISA, $\forall k\in\mathcal{K}_{n}$, generates content from its observation with probability $\beta_{k,n}$ under the following acquisition schemes [11]:
  1. 1.
    Uniform: Content is generated periodically at each time slot, regardless of the prior state of the source and the latest perceived content at the NMAs.
  2. 2.
    Change-aware: Content for an attribute is generated if the state of that attribute changes at a time slot, regardless of the latest perceived content.
  3. 3.
    Semantics-aware: Content generation for an attribute is triggered once there is a discrepancy between the current actual state of the attribute and that of the latest decoded one of the same level at the NMAs.
Then, the ISA follows a self-decision policy, thanks to which it individually decides whether to transmit that content relevant to the $n$-th attribute with probability $\alpha_{k,n}$. Content is assigned a meta-value according to its usefulness in satisfying the subscribed goal [12]. We consider $v_{k,i}^{(n)}$, $\forall k,i,n$, the meta-value for the realization $x_{i}^{(n)}$ at the $k$-th ISA, and assume it is assigned based on a knowledge fusion model, as in [12, Section III-A]. In this work, the self-decision scheme takes the form of a threshold-based decision policy. Thus, the $k$-th ISA speaks up at the $j$-th slot if $v_{k,j}^{(n)}>v_{{\rm th},k}^{(n)}$; otherwise, it remains silent. The threshold $v_{{\rm th},k}^{(n)}$, shows a variable decision-making criterion for the $k$-th ISA observing the $n$-th attribute.
II-B Transmission Success Probability
As observations are noisy, contents generated from the same event via different ISAs can be different, which leads to signal contamination over the wireless network. We consider $h_{k,m}=\bar{h}_{k,m}d_{k,m}^{-\frac{a}{2}}$ the channel between the $k$-th ISA and the $m$-th NMA, represented by an independent and identically distributed (i.i.d.) random variable (r.v.) where $\bar{h}_{k,m}$ is the small-scale fading with its power having a cumulative distribution function (CDF) of $F_{h^{2}}(|\bar{h}_{k,m}|^{2})=1-\operatorname{exp}(-|\bar{h}_{k,m}|^{2})$, $d_{k,m}$ denoting the i.i.d. distance r.v.s, and $72\leq a<$ 7 being the path-loss exponent. Let us consider $\gamma_{\rm th}$ the signal-to-noise ratio (SNR) sensitivity. The probability of successful transmission for content sent from the $k$-th ISA and received at all NMAs is $\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}=\mathrm{Pr}\big{(}\operatorname{SNR}_{k,n}^{(\mathcal{C},\mathcal{I})}>\gamma_{\rm th}\big{)}$. In this regard, the channel drop-off occurs when $\operatorname{SNR}_{k,n}^{(\mathcal{C},\mathcal{I})}\leq\gamma_{\rm th}$, with the probability of $1-\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}$. Here, $\mathcal{C}$ shows the set of collaborating ISAs, and $\mathcal{I}$ is that of interfering ones. After some algebraic manipulations, we reach
$\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}$ $=\sum_{m=1}^{M}\!\Bigg{[}\sum_{k^{\prime}\in\mathcal{C}_{k}}\sum_{k^{\prime\prime}\in\mathcal{I}}\dfrac{\prod_{k^{\prime}_{0}\in\mathcal{C}_{k}}\Lambda_{k^{\prime}_{0},m}^{(n)}\prod_{k^{\prime\prime}_{0}\in\mathcal{I}}\Omega_{k^{\prime\prime}_{0},m}^{(n)}}{\!\left(\Lambda_{k^{\prime},m}^{(n)}\!+\!\Omega_{k^{\prime\prime},m}^{(n)}\right)\!\Lambda_{k^{\prime},m}^{(n)}\Psi_{k,k^{\prime},k^{\prime\prime},m}^{(n)}}$
×$~{}~{}~{}~{}~{}~{}\times\operatorname{exp}\!\left(-\Lambda_{k^{\prime},m}^{(n)}\gamma_{\rm th}\sigma_{m}^{2}\right)\!\Bigg{]}$ (1)
where $\mathcal{C}_{k}=\mathcal{C}\cup\{k\}$, $\Lambda_{k^{\prime},m}^{(n)}=\frac{d_{k^{\prime},m}^{a}}{\rho_{k^{\prime},j}^{(n)}}$, $\Omega_{k^{\prime\prime},m}^{(n)}=\frac{d_{k^{\prime\prime},m}^{a}}{\gamma_{\rm th}\rho_{k^{\prime\prime},j}^{(n)}}$,
$\Psi_{k,k^{\prime},k^{\prime\prime},m}^{(n)}=\!\!\prod_{i\in\mathcal{C}_{k}\setminus\{k^{\prime}\}}\!\!\!\left(\Lambda_{i,m}^{(n)}\!-\!\Lambda_{k^{\prime},m}^{(n)}\right)\!\times\!\!\prod_{j\in\mathcal{I}\setminus\{k^{\prime\prime}\}}\!\!\!\left(\Omega_{j,m}^{(n)}\!-\!\Omega_{k^{\prime\prime},m}^{(n)}\right)\!,$
$\sigma_{m}^{2}$ is the power of the complex additive white Gaussian noise (AWGN) at the $m$-th NMA, and $\rho_{k,j}^{(n)}$ is the transmission power of the $k$-th ISA for the $n$-th attribute at the $j$-th time slot.
III Performance Metrics
We introduce a new metric, named grade of effectiveness (GoE), which measures the effectiveness of the content communicated over the multiple access network to satisfy the subscribed goal. This metric is defined as $\mathcal{G}=\varphi(g(\mathcal{I}))$, where $g:\mathbb{R}^{c}\to\mathbb{R}^{c^{\prime}},c\geq c^{\prime}$, is a (nonlinear) function of $c\in\mathbb{Z}^{+}$ content effectiveness features $\mathcal{I}\in\mathbb{R}^{c}$, and $\varphi:\mathbb{R}^{c^{\prime}}\to\mathbb{R}$ is a context-dependent, cost-aware function [2].
In this letter, we consider the following three effectiveness features: effective discrepancy error (EDE), effective resource consumption (ERC), and effective utility of content (EUoC).
III-A Effective Discrepancy Error (EDE)
A content discrepancy error $E_{t}$ occurs at the $t$-th service period if the actual event $X_{t}$ is not equal to the reconstructed $\widehat{X}_{t}$, i.e., $E_{t}=\mathbbm{1}\{X_{t}\neq\widehat{X}_{t}\}$. The EDE depicts the feature that the effect of losing information is scaled by its usefulness in satisfying the goal. Given this, we model the EDE as follows
${\rm EDE}(E_{t},\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})=\mathcal{F}_{1}\sum_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}P_{{\rm e},n}$ (2)
where $\text{where~{}}\mathcal{F}_{1}$ $=\sum_{j=1}^{{\mid\widehat{\mathcal{A}}_{t}\mid}}\left[\,\sum_{k\in\mathcal{K}_{n}}\alpha_{k,n}v_{k,j}^{(n)}\right]_{n\leftarrow\mathcal{A}_{j}},$
$\boldsymbol{\alpha}=[\alpha_{k,n}]_{K\times N}$, $\mathbf{v}=[v_{k,j}^{(n)}]_{K\times\mid\widehat{\mathcal{A}}_{t}\mid\times N}$, $\mathcal{A}_{j}\subseteq\widehat{\mathcal{A}}_{t}$ shows the attribute queried at the $j$-th time slot, and the notation $n\leftarrow\mathcal{A}_{j}$ implies that $\mathcal{A}_{j}$ is the $n$-th attribute. We also have
$P_{{\rm e},n}=\underset{T\to\infty}{\lim}\frac{1}{T}\sum_{t=1}^{T}\bigg{[}\mathrm{Pr}\!\left({E}_{t,j}^{(n)}=1\right)\!\bigg{]}_{j:\mathcal{A}_{j}\in\widehat{\mathcal{A}}_{t}},$ (3)
which is the expected discrepancy error probability for the $n$-th attribute and is derived according to Proposition 1. In (3), we define ${E}_{t,j}^{(n)}=\mathbbm{1}\{x_{j}^{(n)}\neq\widehat{x}_{j}^{(n)}~{}|~{}t\}$, and $\widehat{x}_{j}^{(n)}$ shows the decoded realization via the NMAs for the $n$-th attribute at the $j$-th slot.
Proposition 1.
The expected discrepancy error probability $P_{{\rm e},n}$ for the $n$-th attribute is derived as
$P_{{\rm e},n}=\dfrac{\mathcal{E}_{n}(I_{n}\!-\!1)p_{n}}{1+\mathcal{E}_{n}\!\left[2(I_{n}\!-\!1)p_{n}-\frac{I_{n}-2}{I_{n}-1}p_{n}-1\right]}$ (4)
where $\mathcal{E}_{n}$ denotes the probability that the $n$-th attribute is not successfully delivered to at least $\lceil\frac{M+1}{2}\rceil$ NMAs, defined as
${\mathcal{E}}_{n}=\sum_{\ell=1}^{2^{|\mathcal{K}_{n}|}}\tilde{\mathcal{E}}_{n,\ell}$ (5)
with ~$\tilde{\mathcal{E}}_{n,\ell}$ given by
~$\tilde{\mathcal{E}}_{n,\ell}=\prod_{r^{\prime}\in\mathcal{R}^{\prime}_{\ell}}(1\!-\!\alpha_{r^{\prime},n})\prod_{r\in\mathcal{R}_{\ell}}\alpha_{r,n}$
×$~{}\times\sum_{\jmath=1}^{2^{|\mathcal{R}_{\ell}|}}\Bigg{[}\prod_{\tau^{\prime}\in\mathcal{T}^{\prime}_{\jmath}}(1\!-\!\chi_{\tau^{\prime},n})\prod_{\tau\in\mathcal{T}_{\jmath}}\chi_{\tau,n}\prod_{v\in\mathcal{V}_{\jmath}}\!\left(1\!-\!\gamma_{v,n}^{(\mathcal{T}_{\jmath},\mathcal{T}_{\jmath}^{\prime})}\right)\!\Bigg{]}$ (6)
where $\chi_{k,n}$ is the probability that the $k$-th ISA’s observation is correct. Also, $\mathcal{R}_{\ell}$, $\mathcal{R}^{\prime}_{\ell}$, $\mathcal{T}_{\jmath}$, and $\mathcal{T}^{\prime}_{\jmath}$ are the sets of ISAs that are active, inactive, active with correct observations, and active with incorrect observations, respectively. Moreover, $\mathcal{V}_{\jmath}$ shows the subset of $\lceil\frac{M}{2}\rceil$ ISAs from $\mathcal{T}_{\jmath}$ with the longest distances from all NMAs if $\mid\!\mathcal{T}_{\jmath}\!\mid\,\geq\lceil\frac{M}{2}\rceil$, $\forall\jmath$; otherwise, $\mathcal{V}_{\jmath}$ is a null set.
Proof:
We use the same approach as in [11, Section 3.2] to derive (4). Considering its definition, the discrepancy error $E_{t,j}^{(n)}$, $\forall t,j,n$, follows a DTMC with two possible states ${\{0,1\}}$ and transition probabilities $\pi^{(n)}_{a,b}=\mathrm{Pr}\big{(}E_{t+1,j^{\prime}}^{(n)}=b~{}|~{}E_{t,j}^{(n)}=a\big{)}$, $\forall a,b\in\{0,1\}$, where $\mathcal{A}_{j}=\mathcal{A}_{j^{\prime}}=n$. Therefore, we can obtain the steady-state error probability $P_{{\rm e},n}$, as follows
$P_{{\rm e},n}=\dfrac{\pi^{(n)}_{0,1}}{1+\pi^{(n)}_{0,1}-\pi^{(n)}_{1,1}}.$ (7)
After computing $\pi^{(n)}_{0,1}$ and $\pi^{(n)}_{1,1}$, we insert them into (7) and obtain (4) after some algebraic manipulations. ∎
III-B Effective Resource Consumption (ERC)
The ERC feature allows us to model the resource consumption for communicating the content as a function of its usefulness. Having time-slotted channels with shared frequency band, power is the main candidate resource to adapt. Therefore, we define the transmission power of the $k$-th ISA at the $j$-th time slot as $\rho_{k,j}^{(n)}=f_{\rho}(v_{k,i}^{(n)})$, in the form of a non-decreasing function $f_{\rho}:\mathbb{R}^{+}_{0}\to\mathbb{R}^{+}_{0}$ of the meta-value assigned to $x_{i}^{(n)}$ at the $j$-th slot. In this regard, we can write
${\rm ERC}(\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})=\mathcal{F}_{2}\prod_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}\mathcal{S}_{n}$ (8)
where $\text{where~{}}\mathcal{F}_{2}=\sum_{j=1}^{{\mid\widehat{\mathcal{A}}_{t}\mid}}\left[\sum_{k\in\mathcal{K}_{n}}h\!\left(\alpha_{k,n}\dfrac{f_{\rho}(v_{k,j}^{(n)})}{v_{k,j}^{(n)}}\right)\right]_{n\leftarrow\mathcal{A}_{j}}\!,$
$h:\mathbb{R}^{+}_{0}\to\mathbb{R}^{+}_{0}$, and $\mathcal{S}_{n}=1-\mathcal{E}_{n}$ is the probability of correctly delivering the $n$-th attribute to the required number of NMAs.
III-C Effective Utility of Content (EUoC)
The EUoC feature shapes the utility (usefulness) of the contents in satisfying the goal once the actual event they are representing is correctly reconstructed at the NMAs. Thereby, the EUoC at each service period depends on the sum value of the communicated attributes at that period, modeled as follows
${\rm{EUoC}}(\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})=\mathcal{F}_{1}\prod_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}\mathcal{S}_{n}.$ (9)
IV Multiple Access Design
We aim to design a multiple access scheme via obtaining the set of optimal activation probabilities $\boldsymbol{\alpha}^{*}=[\alpha_{k,n}^{*}]_{K\times N}$, maximizing the GoE $\mathcal{G}$ in the system based on the goal. Next, we find the threshold criteria for decision-making and propose a self-decision scheme under all content acquisition schemes.
IV-A Problem Formulation
With regard to the definition of $\mathcal{G}$ in Section III, we formulate the following optimization problem
$\mathcal{P}:~{}$ $\underset{\boldsymbol{\alpha},\mathbf{W}}{\text{max}}~{}\mathcal{G}=\varphi\Big{(}\mathbf{W};g_{1}\big{(}I_{1}\big{)},g_{2}\big{(}I_{2}\big{)},...,g_{c}\big{(}I_{c}\big{)}\Big{)}$
[…]
where $0\prec\mathbf{W}=[{w}_{r}]_{c\times 1}\preceq 1$ is a weight vector. We reform the GoE optimization in $\mathcal{P}$ to a problem of minimizing the weighted sum of the EDE and ERC features, subject to guaranteeing that the EUoC exceeds the minimum desirable effectiveness $U_{\rm GoE}$ at each service period. As transmissions are performed independently among the service periods, it is sufficient to analyze the effectiveness problem for an arbitrary $t$-th period. Then we can apply the same access design for other periods, depending on $\widehat{\mathcal{A}}_{t}$, $\forall t$. Finally, the problem becomes
$\widehat{\mathcal{P}}:~{}$ $\underset{\boldsymbol{\alpha},\mathbf{W}}{\text{min}}\left[w_{1}g_{1}\big{(}{\rm EDE}(E_{t},\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})\big{)}\!+\!w_{2}g_{2}\big{(}{\rm ERC}(\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})\big{)}\right]$
[…]
where $g_{r}:\mathbb{R}_{0}^{+}\to\mathbb{R}$, for $r\in\{1,2,3\}$, indicates a positive non-decreasing differentiable function.
IV-B Problem Solution
Inserting (2), (8), and (9) into (IV-A), we define the Lagrange function $\mathcal{L}$ for $\widehat{\mathcal{P}}$, and we can write the Karush-Kuhn-Tucker (KKT) conditions, as follows
$\dfrac{\partial\mathcal{L}}{\partial\alpha_{k,n}}$ $=w_{1}\Bigg{[}\mathcal{F}_{1}\!\left(\dfrac{\partial P_{{\rm e},n}}{\partial\alpha_{k,n}}\right)\!+\!P_{{\rm e},n}\!\left(\dfrac{\partial\mathcal{F}_{1}}{\partial\alpha_{k,n}}\right)\!\Bigg{]}g_{1}^{\prime}\!\!\left(\mathcal{F}_{1}\sum_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}P_{{\rm e},n}\right)$
$~{}~{}~{}+w_{2}\overline{\mathcal{S}}_{n}\Bigg{[}\mathcal{S}_{n}\!\left(\dfrac{\partial\mathcal{F}_{2}}{\partial\alpha_{k,n}}\right)\!+\!\mathcal{F}_{2}\!\left(\dfrac{\partial\mathcal{S}_{n}}{\partial\alpha_{k,n}}\right)\!\Bigg{]}g_{2}^{\prime}\!\left(\mathcal{F}_{2}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)$
$~{}~{}~{}-\eta\overline{\mathcal{S}}_{n}\Bigg{[}\mathcal{S}_{n}\!\left(\dfrac{\partial\mathcal{F}_{1}}{\partial\alpha_{k,n}}\right)\!+\!\mathcal{F}_{1}\!\left(\dfrac{\partial\mathcal{S}_{n}}{\partial\alpha_{k,n}}\right)\!\Bigg{]}g_{3}^{\prime}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)$
$~{}~{}~{}+\lambda_{k,n}=0,\,\forall k,n,$ (12)
where $\overline{\mathcal{S}}_{n}=\!\!\prod_{n^{\prime}\in\widehat{\mathcal{A}}_{t}\setminus{\{n\}}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid-1}}}\!\mathcal{S}_{n^{\prime}}$. Moreover, we have
$\dfrac{\partial\mathcal{L}}{\partial w_{1}}$ $=g_{1}\!\!\left(\mathcal{F}_{1}\sum_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}P_{{\rm e},n}\right)\!+\mu_{0}+\mu_{1}=0,$ (13)
$\dfrac{\partial\mathcal{L}}{\partial w_{2}}$ $=g_{2}\!\left(\mathcal{F}_{2}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)\!+\mu_{0}+\mu_{2}=0.$ (14)
In (IV-B)–(14), $\eta\geq 0$, $\mu_{r}\geq 0$, $\forall r$, and $\lambda_{k,n}\geq 0$, $\forall n\in\widehat{\mathcal{A}}_{t}$, $\forall k\in\mathcal{K}_{n}$, denote Lagrange multipliers.
The complementary slackness conditions are as below
$\mathcal{H}_{1}:\eta\Big{(}g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)\!-U_{\rm GoE}\Big{)}=0,$
$\mathcal{H}_{2}:\mu_{r}\!\left(w_{r}\!-\!1\right)\!$ $=0,\,r=1,2,~{}\,\mathcal{H}_{3}:\lambda_{k,n}\!\left(\alpha_{k,n}\!-\!1\right)\!=0,\,\forall k,n.$
Each of $\mathcal{H}_{1}$ to $\mathcal{H}_{3}$ holds under two conditions: In $\mathcal{H}_{3}$, we consider (i) $\lambda_{k,n}=0$, hence $\alpha_{k,n}<1$; or (ii) $\alpha_{k,n}=1$, thus $\lambda_{k,n}>0$. Under condition (ii), the ISAs are active at all time slots, which is not optimal according to (4) and (5). Thus, we reach $\lambda_{k,n}=0$, $\forall k,n$. In $\mathcal{H}_{2}$, we assume (i) $\mu_{r}=0$, then $w_{r}<1$; or (ii) $w_{r}=1$, hence $\mu_{r}>0$. Condition (ii) leads one of the weighting variables to become zero, which is incorrect referring to $\widehat{\mathcal{P}}$. Thus, we have $\mu_{r}=0$, $\forall r$. Given this, we arrive
$\mathcal{F}_{2}=\dfrac{1}{\mathcal{S}_{n}\overline{\mathcal{S}}_{n}}g_{2}^{-1}\!\!\left[g_{1}\!\!\left(\mathcal{F}_{1}\sum_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}P_{{\rm e},n}\right)\!\right]\!.$ (15)
Also, for $\mathcal{H}_{1}$, we have (i) $\eta=0$, and $g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)>U_{\rm GoE}$; or (ii) $g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)=U_{\rm GoE}$, and $\eta>0$. Applying (15) into (IV-B), condition (i) results in negative probabilities. Thus, condition (ii) is true with $g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)=U_{\rm GoE}$. From (IV-B), we obtain
$P_{{\rm e},n}=\left[\dfrac{g_{1}^{-1}\!\!\left(\dfrac{\eta\mathcal{S}_{n}\overline{\mathcal{S}}_{n}U_{\rm GoE}}{2w_{1}}\right)}{\mathcal{F}_{1}\sum_{n^{\prime}\in\widehat{\mathcal{A}}_{t}\setminus{\{n\}}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid-1}}}P_{{\rm e},n^{\prime}}}\right]^{\!1/2}\!.$ (16)
Putting together (4), (1), and (16), and after some simple numerical and algebraic manipulations based on the forms of $g_{1}$, $h$, and $f_{\rho}$, the optimal $\alpha^{*}_{k,n}$, $\forall k,n$, is derived. We achieve optimal $w_{1}^{*}$ and $w_{2}^{*}$ by exerting $\boldsymbol{\alpha}^{*}$ into the Lagrange function and equalizing the equation $\mathcal{L}(\mathbf{W})=\mathcal{L}(\boldsymbol{\alpha}^{*})$. In this regard, Algorithm 1 is proposed, which provides the optimal values for the optimization variables with certain stopping accuracy and convergence rate of $\mathcal{O}((N_{i}N_{j})^{-1})$, where $N_{i}$ and $N_{j}$ are the maximum numbers of inner and outer loops, respectively.
Input: Known parameters $M$, $L$, $p$, $\widehat{\mathcal{A}}_{t}$, $\mathcal{K}_{n}$, $v_{k,j}^{(n)}$, $\forall j,k$, $\chi_{k,n}$, $\gamma^{(\mathcal{C},\mathcal{I})}_{k,n}$, $\forall n,k$, $U_{\rm GoE}$, $\boldsymbol{\varepsilon_{1}}$, and $\boldsymbol{\varepsilon_{2}}$. Initial parameters $\eta^{(0)}\to 0$, $\boldsymbol{\alpha}^{(0)}=\mathbf{0}_{K\times|\widehat{\mathcal{A}}_{t}|}$, and $\mathbf{W}=[0.5,0.5]^{T}$. Forms of $g_{r}$, $\forall r$, $h$, and $f_{\rho}$.
Output: Optimal parameters $\alpha^{*}_{k,n}$, $\forall k,n$, $w_{1}^{*}$, and $w_{2}^{*}$.
1 Iteration $j$: $\triangleright$ Outer loop
2 if conditions $\mathcal{H}_{4}$ to $\mathcal{H}_{6}$ are met. then goto 3. Iteration $i$: $\triangleright$ Inner loop
3 Consider $n=\langle\widehat{\mathcal{A}}_{t}(c)\rangle_{c=1}$ and $k=\langle\mathcal{K}_{n}(c)\rangle_{c=1}$.
4 Update $\overline{\mathcal{S}}_{n}^{(i)}$, $\mathcal{F}_{1}^{(i)}$, $\mathcal{F}_{2}^{(i)}\!(\boldsymbol{\alpha})$, and $P^{(l,i)}_{{\rm e},n^{\prime}}$, $\forall n^{\prime}\in\widehat{\mathcal{A}}_{t}\setminus\{n\}$.
5 Derive $\alpha_{k,n}^{*}$ from (4), (5), (1), (16), and given $\alpha_{k^{\prime},n}$, $\forall n,k^{\prime}\in\mathcal{K}_{n}\setminus\{k\}$; update and save $\boldsymbol{\alpha}^{(i)}$.
6 goto 4 and repeat for $\langle\widehat{\mathcal{A}}_{t}(c)\rangle_{c=2}^{|\widehat{\mathcal{A}}_{t}|}$ and $\langle\mathcal{K}_{n}(c)\rangle_{c=2}^{|\mathcal{K}_{n}|}$.
7 if $\Big{|}\boldsymbol{\alpha}^{(i)}\!-\!\boldsymbol{\alpha}^{(i-1)}\Big{|}\!>\!\boldsymbol{\varepsilon_{1}}$ then set $i\!=\!i\!+\!1$, and goto 3. Compute $w_{1}^{(j)}$ and $w_{2}^{(j)}$; save $\mathbf{W}^{(j)}$.
if $\Big{|}\mathbf{W}^{(j)}\!-\!\mathbf{W}^{(j-1)}\Big{|}\!>\!\boldsymbol{\varepsilon_{2}}$, $\mathbf{W}^{(j)}\preceq 0$, $\mathbf{W}^{(j)}\succeq 1$, or $\exists_{k,n}\alpha_{k,n}^{(i)}<0$ then step up $\eta^{(j)}$, $j\!=\!j\!+\!1$, and goto 1. Return $\boldsymbol{\alpha}^{*}=\boldsymbol{\alpha}^{(i)}$ and $\mathbf{W}^{*}=\mathbf{W}^{(j)}$.
Algorithm 1 Solution for deriving $\boldsymbol{\alpha}^{*}$ and $\mathbf{W}^{*}$ at the $t$-th service period.
Remark 1.
The convexity of ${\mathcal{L}}$ relies on the forms of $g_{1}$, $g_{2}$, and $g_{3}$. For convex $g_{1}$ and $g_{2}$ but concave $g_{3}$, the dual problem is definitely convex with the following conditions
$\mathcal{H}_{4}$ $:\dfrac{\partial\mathcal{F}_{1}}{\partial\alpha_{k,n}}\dfrac{\partial P_{{\rm e},n}}{\partial\alpha_{k,n}}+\dfrac{\mathcal{F}_{1}}{2}\dfrac{\partial^{2}P_{{\rm e},n}}{\partial\alpha_{k,n}^{2}}\geq 0,$
$\mathcal{H}_{5}$ $:\dfrac{\partial\mathcal{S}_{n}}{\partial\alpha_{k,n}}\dfrac{\partial\mathcal{F}_{2}}{\partial\alpha_{k,n}}+\dfrac{\mathcal{S}_{n}}{2}\dfrac{\partial^{2}\mathcal{F}_{2}}{\partial\alpha_{k,n}^{2}}\geq 0.$
For all being convex, ${\mathcal{L}}$ is convex if $\mathcal{H}_{4}$, $\mathcal{H}_{5}$, and $\mathcal{H}_{6}$ hold, with
$\mathcal{H}_{6}$ $:\eta\left(\mathcal{F}_{1}\dfrac{\partial\mathcal{S}_{n}}{\partial\alpha_{k,n}}\!+\!\mathcal{S}_{n}\dfrac{\partial\mathcal{F}_{1}}{\partial\alpha_{k,n}}\right)g_{3}^{\prime\prime}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)\!$
$~{}~{}~{}~{}~{}~{}\leq w_{2}\left(\mathcal{F}_{2}\dfrac{\partial\mathcal{S}_{n}}{\partial\alpha_{k,n}}\!+\!\mathcal{S}_{n}\dfrac{\partial\mathcal{F}_{2}}{\partial\alpha_{k,n}}\right)g_{2}^{\prime\prime}\!\left(\mathcal{F}_{2}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)\!.$ (17)
The reverse of Remark 1 is not necessarily true. Exerting (8), (15), and Remark 1, optimal forms of $h$ and $f_{\rho}$ are found.
IV-C Analysis for Large $I_{n}$
We consider the attribute $\mathcal{X}_{n}$, $\forall n$, has large enough state space, i.e., $I_{n}\gg 1$. In this regime, we achieve $\underset{L\to\infty}{\lim}P_{{\rm e},n}=\frac{1}{2}$ from (4). Thus, using (16), $\alpha^{*}_{k,n}$, $\forall k,n$, is derived as follows
$\alpha_{k,n}^{*}=\dfrac{\eta\overline{\mathcal{S}}_{n}U_{\rm GoE}\Big{(}1\!-\widehat{\mathcal{E}}_{n}^{(k)}\Big{)}-2w_{1}g_{1}\!\!\left(\dfrac{\mathcal{F}_{1}}{2^{(|\widehat{\mathcal{A}}_{t}|+1)}}\right)}{\eta\overline{\mathcal{S}}_{n}U_{\rm GoE}\Big{(}{\mathcal{E}}_{n}^{(k)}\!-\widehat{\mathcal{E}}_{n}^{(k)}\Big{)}\!}$ (18)
where we rewrite (1) and define ${\mathcal{E}}_{n}^{(k)}$ and ^$\widehat{\mathcal{E}}_{n}^{(k)}$ as below
${\mathcal{E}}_{n}^{(k)}=\dfrac{\sum_{\ell:k\in\mathcal{R}_{\ell}}\tilde{\mathcal{E}}_{n,\ell}}{\alpha_{k,n}},~{}\,\widehat{\mathcal{E}}_{n}^{(k)}=\dfrac{\sum_{\ell:k\in\mathcal{R}^{\prime}_{\ell}}\tilde{\mathcal{E}}_{n,\ell}}{1-\alpha_{k,n}}.$
From (18), we infer that $\alpha_{k,n}^{*}$ increases by (i) decreasing the average usefulness of contents to tackle the high contamination of useful contents over the channels, (ii) enlarging the set of queried attributes, and (iii) increasing the $U_{\rm GoE}$ constraint.
IV-D Decision Making Criterion
To obtain the threshold criterion $v_{{\rm th},k}^{(n)}$, $\forall k$, $\forall n\in\mathcal{A}_{k}$, for the self-decision scheme proposed in II-A, the $k$-th ISA needs to timely compute $\alpha_{k,n}^{*}$, $\forall k,n$, according to Section IV-B. To this end, the ISA measures its distance from all NMAs thanks to the received query signals and has knowledge of $\mathcal{K}_{n}$, the other ISAs’ location distribution, content usefulness distributions, and observation accuracy. We derive the following upper bound for $\gamma_{k,n}$ in (II-B) by applying the Markov’s inequality:
$\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}\leq$
$~{}~{}~{}\sum_{m=1}^{M}\underset{\mathcal{C}_{k}}{\mathbb{E}}\!\left\{\sum_{k^{\prime}\in\mathcal{C}_{k}}\sum_{k^{\prime\prime}\in\mathcal{I}}\dfrac{\prod_{k^{\prime}_{0}\in\mathcal{C}_{k}}\Lambda_{k^{\prime}_{0},m}^{(n)}\prod_{k^{\prime\prime}_{0}\in\mathcal{I}}\Omega_{k^{\prime\prime}_{0},m}^{(n)}}{\!\left(\Lambda_{k^{\prime},m}^{(n)}\!+\!\Omega_{k^{\prime\prime},m}^{(n)}\right)\!\!\left(\Lambda_{k^{\prime},m}^{(n)}\right)^{\!3}}\,\Omega_{k^{\prime\prime},m}^{(n)}\right\}\!.$ (19)
After equalizing $\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}$ to its upper bound and using Algorithm 1, $\alpha_{k,n}^{*}$ is derived for the $k$-th ISA and $n$-th attribute at any arbitrary service period. Eventually, we arrive
$v_{{\rm th},k}^{(n)}=F_{V}^{-1}\!\!\left(1\!-\!\operatorname{min}\!\left\{1,\dfrac{\alpha_{k,n}^{*}}{\operatorname{max}\!\big{\{}0,\beta_{k,n}\big{\}}}\right\}\right)$ (20)
where $F_{V}(v_{k,j}^{(n)})$ indicates the CDF of the content meta-value. Also, $\beta_{k,n}$ shows the scheme-dependent probability that the $k$-th ISA generates contents for the $n$-th attribute, given by
$\beta_{k,n}=\begin{cases}1;\hskip 116.6563pt\text{uniform}\\ (I_{n}\!-\!1)p_{n};\hskip 79.66771pt\text{change-aware}\\ P_{{\rm e},n}(1\!-\!I_{n}p_{n})+(I_{n}\!-\!1)p_{n};\hskip 7.39772pt\text{semantics-aware}\end{cases}$ (21)
if $k\in\mathcal{K}_{n}$; otherwise, $\beta_{k,n}=0$.
V Simulation Results
We investigate the performance of the proposed self-decision scheme in the system, satisfying the goal with $U_{\rm GoE}=1$. By default, we set $K=2$, $M=1$, $I_{n}=5$, $\forall n$, $p=0.2$ (fast-changing regime), and $\chi_{k,n}=0.8$, $\forall k,n$. ISAs are distributed uniformly in an area of $50\times 50\,[\text{m}^{2}]$, and NMAs are located in a formation to cover all ISAs. The assignment of a meta-value to each content follows a Gaussian distribution with mean $0.5$ and variance $0.2$. Besides, $P_{X}(\cdot)$ denotes the same distribution, but with mean 00 and variance $5$. Moreover, $g_{1}$, $g_{2}$, and $g_{3}$ have exponential forms, while $h$ and $f_{\rho}$ are linear. Also, the path-loss exponent $a=3.8$, $\sigma_{m}^{2}=-40\,[\text{dBm}]$, and $\gamma_{\rm th}=25\,[\text{dB}]$.
Fig. 2 illustrates the system’s effectiveness performance using the obtained optimal activation probabilities and weight factors of the considered features. The semantics-aware scheme empowered with the self-decision scheme represents the highest GoE for communicated contents from different viewpoints. This performance comes with a smaller discrepancy error, higher useful content acquisition, and fewer drop-offs compared to the other schemes. For instance, in Fig. 2 (a), the semantics-aware scheme offers on average $2\%$ and $2.19\%$ higher GoE with $25.13\%$ and $6.09\%$ lower drop-offs of contents than those of the uniform and change-aware schemes, respectively. Besides, the self-decision scheme adds $21.38\%$, $20.6\%$, and $27.33\%$ to the GoE and decreases the transmission rate by $42.2\%$, $67.21\%$, and $48.3\%$ for the semantics-aware, uniform, and change-aware schemes, respectively, compared to the case with no decision-making, where $\alpha_{k,n}=\beta_{k,n}$, $\forall k,n$. In a nutshell, the joint self-decision scheme and the semantics-aware scheme provide on average up to $29.52\%$ higher GoE, $25.13\%$ fewer drop-offs, and $67.21\%$ lower transmission rate.
[…]
[scale=0.215]Results/Dimension_NEW.eps
[…]
[scale=0.215]Results/NMA_ISASize_NEW.eps
[…]
[scale=0.215]Results/AttributeSize_NEW.eps
[…]
[scale=0.215]Results/StateSize_NEW.eps
Figure 2: The interplay between GoE and (a) the coverage area dimension, (b) the number of ISAs and NMAs, (c) the number of queried attributes, and (d) the size of attribute state space.
Also, as in Fig. 2 (a), widening the coverage area reduces the GoE under a fixed number of agents since the transmission success probability dwindles with distance. Moreover, Fig. 2 (b) shows that increasing $K$ and/or $M$ enhances the GoE. Although a larger $K$ boosts the coverage density, hence the success probability, it gradually raises the number of interfering agents. Therefore, the offered GoE is saturated for large enough $K$. According to Fig. 2 (c), the GoE decreases if the NMAs require more attribute levels, since discrepancy errors are accumulated at each service period. In Fig. 2 (d), increasing $I_{n}$, $\forall n$, brings in higher EDE and lower ERC with respect to the EUoC, while counterbalancing each other’s impacts for larger $I_{n}$. For the change-aware scheme with $I_{n}<6$, the ERC is larger than the other features, as the system works in a fast-changing regime.
VI Conclusion
We developed a self-decision goal-oriented multiple access scheme for networked intelligent systems. Aiming to maximize the GoE, we obtained the optimal activation probabilities and threshold criteria for decision-making under different content acquisition schemes. Our results showed that our proposal could provide on average up to $29.52\%$ higher GoE, $25.13\%$ fewer drop-offs, and $67.21\%$ lower transmission rate.